OpenStudy (desmarie):

A rocket blasts off. In 10.0 seconds it is at 10,000 ft, traveling at 3600 mph. Assuming the direction is up, calculate the acceleration. (Hint: the rocket is not under constant acceleration). 5280 ft/s2 528 ft/s2 100 ft/s2 200. ft/s2

OpenStudy (desmarie):

@zepdrix @FaiqRaees @sweetburger

OpenStudy (sweetburger):

First we need to convert 3600mph to ft/s \[3600\frac{ miles }{ hour }\times \frac{ 1 hour }{ 60 minutes }\times \frac{ 1 minute }{ 60 seconds }\times \frac{ 5280ft }{ 1 mile }\]

OpenStudy (sweetburger):

brb

OpenStudy (desmarie):

I had the answer as 5280 but it was wrong -.-

OpenStudy (desmarie):

@sweetburger

OpenStudy (desmarie):

@welshfella

OpenStudy (johnweldon1993):

That calculation above was merely to convert miles per hour to ft per s...which is needed here as we need to keep units consistent The next step would have been to recognize that \[\large a = \frac{\Delta v}{\Delta t} = \frac{\text{Final Velocity - Initial Velocity}}{\text{Final time - Initial Time}} = \frac{5280[\frac{ft}{s}] - 0[\frac{ft}{s}]}{10[s] - 0[s]}\] \[\large a = \frac{5280 [\frac{ft}{s}]}{10[s]} = ?\]

OpenStudy (desmarie):

528 ft/s2